Exercise 2-4 Computing Total Job Costs and Unit Product Costs Using Multiple Predetermined Overhead Rates [LO2-4]

Fickel Company has two manufacturing departments—Assembly and Testing & Packaging. The predetermined overhead rates in Assembly and Testing & Packaging are $25.00 per direct labor-hour and $21.00 per direct labor-hour, respectively. The company’s direct labor wage rate is $27.00 per hour. The following information pertains to Job N-60:

Assembly Testing & Packaging
Direct materials $ 405 $ 51
Direct labor $ 297 $ 54

Required:

1. What is the total manufacturing cost assigned to Job N-60? (Do not round intermediate calculations.)

2. If Job N-60 consists of 10 units, what is the unit product cost for this job? (Do not round intermediate calculations. Round your answer to 2 decimal places.)

1. Total manufacturing cost
2. Unit product cost per unit

Garrison 16e Rechecks 2017-08-08

Garrison_16e_Rechecks_2019_10_02
Answer #1
Answer

Direct labor hours:

Assembly = 297/27= 11
Testing & Packaging = 54/27=2
Requirement- 1
Direct materials (405+51) 456
Direct labor (297+54) 351
Overhead applied:
Assembly   (11*25) 275
Testing & Packaging (2*21) 42
Total manufacturing cost 1124
Requirement- 2
Total manufacturing cost 1124
Divide by Units 10
Unit product cost   ( 1124/10) 112.4 per unit

Please like